OpenStudy (anonymous):

I've got to solve this initial value problem dy/dx = 2x / 3y^2-12 with y(-sqrt(65)) = 0

OpenStudy (anonymous):

\[dy/dx = (2x)/(3y ^{2}-12)\]

OpenStudy (anonymous):

Need to find C and then put it in implicit form

OpenStudy (nikvist):

\[\frac{dy}{dx}=\frac{2x}{3y^2-12}\] \[(3y^2-12)dy=2xdx\] \[y^3-12y=x^2+C\] \[y(-\sqrt{65})=0\quad\Rightarrow\quad C=-65\] \[y^3-12y=x^2-65\]